Find the center of mass of the region bounded by y=(x-2)^2 and y=4.



1
Expert's answer
2022-05-30T16:58:34-0400
"(x-2)^2=4"

"x_1=0, x_2=4"


"A=\displaystyle\int_{0}^{4}(4-(x-2)^2)dx=[4x-\dfrac{(x-2)^3}{3}]\begin{matrix}
4 \\
 0
\end{matrix}"

"=16-\dfrac{8}{3}-(0+\dfrac{8}{3})=\dfrac{32}{3}"

"\bar{x}=\dfrac{1}{A}\displaystyle\int_{0}^{4}x(4-(x-2)^2)dx"

"=\dfrac{3}{32}\displaystyle\int_{0}^{4}(-x^3+4x^2)dx=\dfrac{3}{32}[-\dfrac{x^4}{4}+\dfrac{4x^3}{3}]\begin{matrix}
4 \\
 0
\end{matrix}"

"=\dfrac{3}{32}(-64+\dfrac{256}{3})=2"


"\bar{y}=\dfrac{1}{A}\displaystyle\int_{0}^{4}\dfrac{1}{2}((4-(x-2)^2))^2dx"

"=\dfrac{3}{64}\displaystyle\int_{0}^{4}(x^4-8x^3+16x^2)dx"

"=\dfrac{3}{64}[\dfrac{x^5}{5}-2x^4+\dfrac{16x^3}{3}]\begin{matrix}
4 \\
 0
\end{matrix}"

"=\dfrac{3}{64}(\dfrac{1024}{5}-512+\dfrac{1024}{3})=\dfrac{8}{5}"

"(2, \dfrac{8}{5})"